Assume 5 ft3/s of Hydrogen at 15 psia and 80 oF are compressed during an irreversible adiabatic  process to 90 psia and 522 oF. For a non-flow process determine a) k, b) V2, c) W, d) ΔU and e) ΔS.


1
Expert's answer
2021-04-21T07:04:08-0400

a. "\frac{T_2}{T_1}=(\frac{P_2}{P_1})^{\frac{k-1}{k}}"

"\frac{545.372}{299.817}=(\frac{90}{15})^{\frac{k-1}{k}}"

"\ln{1.819}=({\frac{k-1}{k}} )\ln{6}"

"1.1935k=1.79"

"k=1.5"


b. "P_1V_1^k=P_2V_2^k"

"15(5)^{1.5}=90(V_2)^{1.5}"

"V_2=1.5142 In^3\/s"


c. "\triangle w = \frac{P_1V_1-P_2V_2}{k-1}= \frac{15 \times 1728 \times 5 - 90 \times 1782 \times 1.5142}{1.5-1}=-211776.76 lb-inch/s"


d. "\triangle \theta = \triangle w +\triangle u"

But "\triangle \theta =0"

So, "\triangle w = -\triangle u"

"\triangle u=211776.76 lb-inch/s"


e. "\triangle s=\frac{Q}{T}=\frac{-211776.76}{245.555}=-862.44124 lb-inch /sK"